OpenStudy (anonymous):

integrate∫sin 5x cos 2x dx. how's the step to do it.

OpenStudy (mathmagician):

Its easy, if you know, that \[\sin(\alpha)*\cos(\beta)=(\sin(\alpha+\beta)+\sin(\alpha-\beta))/2\]. Then you get \[\int\limits_{}^{}(\sin(7x)+\sin(3x))dx\]. Hope, that helped :)